OpenStudy (anonymous):

Medal Find three consecutive odd integers whose sum is 75. Write your answers in ascending order.

OpenStudy (freckles):

let n be an an odd integer the very next odd integer would be n+2 since n+1 is even and the next one after that would be n+4 since n+3 is even

OpenStudy (freckles):

so let n,n+2, and n+4 be your three odd conseutive integers

OpenStudy (freckles):

add them up and set it's sum to 75

OpenStudy (anonymous):

what are the odd integers

OpenStudy (freckles):

n,n+2, and n+4

OpenStudy (anonymous):

isnt 2 an even number?

OpenStudy (freckles):

2 is an even number yes but what does that have to do with our numbers n,n+2,and n+4 where n is odd

OpenStudy (anonymous):

okay thank you

OpenStudy (freckles):

n+(n+2)+(n+4)=75 solve for n

OpenStudy (freckles):

then find n+2 then find n+4

OpenStudy (anonymous):

i dont understand how we solve

OpenStudy (freckles):

do you know how to combine like terms yet?

OpenStudy (freckles):

n+n+n=? 2+4=?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

6

OpenStudy (freckles):

and n+n+n=?

OpenStudy (anonymous):

3n

OpenStudy (freckles):

n+(n+2)+(n+4)=75 (n+n+n)+(2+4)=75 3n + 6 =75 3n+6=75 can you finish solving? are you haven't solve linear equations yet?

OpenStudy (anonymous):

3n = 69

OpenStudy (anonymous):

23

OpenStudy (freckles):

so n=23 so n+2 is?

OpenStudy (anonymous):

25

OpenStudy (freckles):

and n+4 is?

OpenStudy (anonymous):

27

OpenStudy (anonymous):

23,25,27

OpenStudy (freckles):

and can't you agree that 23,25,27 are odd consecutive integers whose sum is 75?

OpenStudy (anonymous):

yes i agree

OpenStudy (anonymous):

thank you soo much

OpenStudy (freckles):

np